OpenStudy (anonymous):

Consider the sequence formed by the digits of the concatenation of the positive integers, listed in order: 123456789101112131415161718192021... What is the 2011th digit of this sequence?

OpenStudy (anonymous):

@Agent47

OpenStudy (agent47):

what's concatenation? lol

OpenStudy (anonymous):

THis is even more lol... No idea. But from a brief google search, concatenation: the joining of two numbers by their numerals. Like in this one, the numbers are going in order.

OpenStudy (kinggeorge):

Concatenation is basically just putting the two things next to each other. So 25 concatenated with 347 is merely 25347.

OpenStudy (agent47):

Ohhhh

OpenStudy (kinggeorge):

Do you know what the solution is? I would like to check my answer.

OpenStudy (kinggeorge):

For reference, I believe it should be 3.

OpenStudy (anonymous):

The solution is 7

OpenStudy (kinggeorge):

Close...I was one off. (with how I was counting) Let me see where I overcounted.

OpenStudy (anonymous):

The first 7 of 707

OpenStudy (agent47):

mann KG's a genius.

OpenStudy (kinggeorge):

Ah right, I was doing things a bit wrong. It should be 7. Specifically the first 7 of 707. Give me a minute, and I shall explain.

OpenStudy (anonymous):

Wait, I think I got it. Can I post my answer and see if it's right?

OpenStudy (kinggeorge):

sure.

OpenStudy (anonymous):

Well, there are 9 single digit numbers. Then, there are 90 two digit numbers. That gives you 189 digits for 1-99. I have to get to the 2011 digit, so 2001 - 189 = 1822 and the remaining possible digits up to 2011 is 3 digit numbers. 1822/3 = 607 with a remainder of 1, so 607 + 99 = 706 which when adding 1 more digit is 7 of 707 because of the remainder. Does that make sense?

OpenStudy (anonymous):

I meant 2011 near the beginning 2011 - 189

OpenStudy (kinggeorge):

That's exactly what I was doing. Here's the explanation in my words. First, note that you have 9 single digit numbers. Then, you have concatenated 90 two digit numbers. Thus, you have a 9+2(90)=189 digits so far. We must count exactly \(2011-189=1822\) more digits. Since all the next numbers will be 3 digit numbers, we divide by 3. \[1822/3=607.3333333...\]Now, here's the tricky part. The first of these is 100, so we add 607 to 100 to get 707. However, we must also subtract 1 since we're starting at 100, and not 101. Hence, it is the digit immediately after we see \(706\). The next digit is the first in \(707\).

OpenStudy (anonymous):

Alright. Thanks!

OpenStudy (kinggeorge):

You're welcome.